really stuck on this equation with three variables.
3x=-12
2x-y=3z=-1
3x+4-z=-7

Answer by ewatrrr(24785)

 
Hi,
3x=-12 0r x = -4
2x-y+3z=-1
3x+4-z=-7
-8-y+3x = -1 |Substituting x = -4
-12-y -z = -1
-y + 3z = 7
-y - z = 11
-y + 3z = 7
y + z = -11 |Multiplying 2nd EQ by -1
4z = -4
z = -1 and y = -10